OpenStudy (anonymous):

Lost in step 3...? 2. Apply the formula of a right circular cylinder (V = r2h) to find the volume of the object. 3. Rewrite the formula using the variable x for the radius. Substitute the value of the volume found in step 2 for V and express the height of the object in terms of x plus or minus a constant. Diameter: 3in Height: 4.5in Volume: ≈31.8in

OpenStudy (asnaseer):

what this is asking you to do first is to substitute r=x into the equation:\[V = \pi r^2h\]

OpenStudy (asnaseer):

that will give you V interms of x and h

OpenStudy (asnaseer):

then, in step 1 where you found the volume, use that as the value for V and use the value 4.5 for h (the height)

OpenStudy (anonymous):

31.8=(pi)x^2(4.5) ?

OpenStudy (asnaseer):

you should then be able to rearrange the equation to get an expression for x.

OpenStudy (asnaseer):

yes - that looks correct. now rearrange it so that you have x on the left hand side. i.e. you should end up with something like: x = ...

OpenStudy (anonymous):

But what does it mean by in terms of x plus or minus a constant?

OpenStudy (asnaseer):

sorry - I misread the question. you are supposed to find and expression for the height (h) in terms of x.

OpenStudy (asnaseer):

so you have:\[31.8=\pi x^2h\]now rearrange this to get the height (h) on the left hand side.

OpenStudy (asnaseer):

"plus or minus a constant" usually means something like:\[h = 55x^2+12\]or:\[h=55x^2-12\]i.e. the expression for h contains \(x^2\) plus/minus some constant. but, in this case, there should be no constant to add or subtract.

OpenStudy (asnaseer):

BTW: the expressions I showed above are just /examples/ and not the actual answer.

OpenStudy (anonymous):

31.8=(pi)x^2(h) -h+31.8=(pi)x^2 -h=(pi)x^2-31.8 h=-(pi)-^2+31.8

OpenStudy (asnaseer):

lets take it a step at a time...

OpenStudy (asnaseer):

we have:\[31.8=\pi x^2h\]so first lets swap the left and right hand sides as we want h to appear on the left hand side. this gives:\[\pi x^2h=31.8\]make sense so far?

OpenStudy (asnaseer):

think of actual examples, e.g.: 1 + 2 = 3 therefore: 3 = 1 + 2 so swapping left and right hand sides is a legal operation.

OpenStudy (asnaseer):

ok so far?

OpenStudy (asnaseer):

please let me know if this makes sense so far or whether I need to explain more before moving on?